OpenStudy (anonymous):

Which of the follow is pi/3 the reference angle for? (Can choose more than one) (Will give medal & fan) (posting options in comments)

OpenStudy (anonymous):

\[\frac{ 2\pi }{ 3 }\]\[\frac{ 15\pi }{ 3 }\]\[\frac{ 7\pi }{ 3 }\]\[\frac{ 19\pi }{ 3 }\]And if whoever answers this could explain why the answer(s) is/are what it/they is/are, that'd be cool.

OpenStudy (anonymous):

@phi ?

OpenStudy (phi):

Any angle outside of the range 0 to pi/2 (0 to 90º) can be reduced to the reference angle. I would try to plot each angle to see where it is For example the first angle 2pi/3